OpenStudy (anonymous):

9x +3y = -45 2x - 3y = -21

OpenStudy (anonymous):

To solve this you need to first get one of the variables (x or y) by itself on of the sides. Hint: 9x + 3y = -45 simplifies to 3x +y = -15

OpenStudy (radar):

Add the two equations in order to eliminate the y variable: 9x+3y = -45 2x-3y = -21 -------------- add 11x=-66 x=-6 The problem is over half done.

OpenStudy (radar):

Now substitute this newly found value of x in the second equation 2(-6) -3y = -21 -12 - 3y =-21 -3y= -21 +12 -3y = -9 y = 3 Now the problem is solved, x=-6 y=3